OpenStudy (anonymous):

Solve: 4(x + 1) = −4

OpenStudy (anonymous):

Show your steps and we can help you figure it out.

OpenStudy (anonymous):

x = -2 just plug it in to check :)

OpenStudy (anonymous):

Thats the problem i dont know how to start

OpenStudy (radar):

By inspection, I can see x=-2, but lets work it out. Clear the parenthesis: 4x+4=-4, subtract 4 from both sides. 4x=-8 divide both sides by 4 x=-2 yup, it is what it is.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Distributive_property The distributive property states: X*(Y+Z) = (X*Y)+(X*Z)

OpenStudy (anonymous):

okay look you just have to find x so a simple but long way is plugging in numbers but that could take a while so instead you use the distributive property 4x_4=-4 4x+8=0 and so x=-2

OpenStudy (anonymous):

start by distributing the 4 so it would be 4x+4=-4 now you can solve the problem

OpenStudy (anonymous):

ty